OpenStudy (anonymous):

PLEASE I NEED HELP! A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9 (x + 1)2 + (y - 2)2 = 9

OpenStudy (princeharryyy):

(x + 1)2 + (y - 2)2 = 9 general equation>> (x -x1)2 + (y - y1)2 = r^2 where (x1,y1) is the center of the circle

OpenStudy (anonymous):

sorry for the late reply hope youre still there @princeharryyy

OpenStudy (anonymous):

but i have no clue how to work this can you do it step by step with me

OpenStudy (princeharryyy):

you are given x1 = -1, y1 =2 and r = 3 just put it in, in the equation.

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

@princeharryyy

OpenStudy (princeharryyy):

D

OpenStudy (anonymous):

Thank YOU!!! @princeharryyy

OpenStudy (princeharryyy):

:)